OpenStudy (anonymous):

I need help finding the minimum and maximum absolute extreme values, could someone explain to me how to find them using Lagrange???

OpenStudy (anonymous):

a have tons of problems to solve but I really don't get it how to solve them

OpenStudy (goformit100):

post it

OpenStudy (anonymous):

\[f \left( x,y \right)=\left( 4x-2y ^{2} \right)\cos y \]and the restrictions are \[1\le x \le 3,- \Pi/4\le y \le \Pi/4\]

OpenStudy (turingtest):

I don't understand your restriction; aren't you given a function as a restriction g(x,y) ?

OpenStudy (turingtest):

in general you want to find\[f(x,y)=\lambda\nabla g(x,y)\]where\[g(x,y)=k\]is the constraint

OpenStudy (anonymous):

but I do not have a funtion I have the restrictions in the way I wrote it maybe that's why I don't get it

OpenStudy (anonymous):